### v1

# Evidence for simple “arrow of time functions” in closed chaotic quantum systems

Abstract
Through an explicit construction, we assign to any infinite temperature autocorrelation function $C(t)$ a set of functions $\alpha^{n}(t)$. The construction of $\alpha^{n}(t)$ from $C(t)$ requires the first $2n$ temporal derivatives of $C(t)$ at times 00 and $t$. Our focus is on $\alpha^{n}(t)$ that (almost) monotonously decrease, we call these “arrows of time functions” (AOTFs). For autocorrelation functions of few body observables we numerically observe the following: An AOTF featuring a low $n$ may always be found unless the the system is in or close to a nonchaotic regime with respect to a variation of some system parameter. All $\alpha^{n}(t)$ put upper bounds to the respective autocorrelation functions, i.e. $\alpha^{n}(t)\geq C^{2}(t)$. Thus the implication of the existence of an AOTF is comparable to that of the H-Theorem, as it indicates a directed approach to equilibrium. We furthermore argue that our numerical finding may to some extent be traced back to the operator growth hypothesis. This argument is laid out in the framework of the so-called recursion method.
I Introduction
Is there a tendency in closed, nondriven quantum systems to approach equilibrium in a monotonous fashion? And if so, how would one even pinpoint all the somewhat vague notions in this question? Obviously this question is related to the question for the so called “arrow of time”. A full fledged overview over all approaches to this subject, even restricted to the quantum realm, is clearly beyond the possibilities of the paper at hand. Thus, below we only very briefly introduce some concepts that appear to be in the vicinity/form the background of what shall be mainly presented later in the present paper. Readers well acquainted with these subjects may just skip this part.
arrow of time in closed quantum systems: While the full von Neumann entropy in closed system can never approach equilibrium (i.e. increase), other quantities like expectation values of certain observables, reduced density matrices, etc. may do so, in the sense that large deviations from equilibrium become extremely rare, as captured by the term “equilibration on average”. Under rather mild conditions such a scenario is even likely to emerge, as the milestone concepts of typicality and the eigenstate thermalization hypothesis suggest [1, 2, 3]. The caveat of “equilibration on average” with respect to an arrow of time is, that it only establishes the low relative frequency of deviations from equilibrium, but no temporal order for these rare events emerges. The deviations could e.g. persist for a very long period in the beginning if they subsequently vanish for an even larger period, or (re)appear after a very long period of (apparent) closeness to equilibrium. For an illustration of the latter see Ref. [4], for an attempt to deal with the former see Ref. [5].
arrow of time in measured quantum systems: The von Neumann entropy is nondecreasing under projective measurements (if the measurement result is not read out) [6]. Since it is invariant under unitary evolution, it cannot decrease in a sequence of intermittent unitary evolutions and measurements. A continuous version of this scenario is often captured by means of an open system description with a dephasing model [7]. In such descriptions the Kullback-Leibler divergence of the actual state and a stationary state is nonincreasing [8]. Hence, for repeatedly measured systems this Kullback-Leibler divergence may be considered as a nonincreasing distance to equilibrium and thus constitutes an arrow of time. However, the paradigm that only the external measurement drives systems to equilibrium is at odds with a number of recent experiments (e.g., on cold atoms see [9]) and the above concept of equilibration in closed quantum systems.
arrow of time, observational entropy and consistent histories: This approach may be regarded as an synthesis of the open and the closed system concepts. It is based on the observation that some dynamics of a (typically more complex) closed system may appear as if the latter was subject to external measurements. In this case so called “consistent histories” emerge. Whether or not this is the case could in principle be answered by computing a so called “decoherence functional” [10], in practice this may be hard. However, if consistent histories emerge, the results from the open quantum quantum system carry over and a suitably defined analog of the von Neumann entropy is nonincreasing [11]. It has also been pointed out that in this case an even more accessible quantity called “observational entropy” [12, 13, 14] is nonincreasing. While these are promising perspectives, it is unclear whether consistent histories always emerge when the notion of an arrow of time appears applicable.
In the paper at hand we are working within the closed system framework. However, we aim at going beyond the common lore of “scarceness of excursions from equilibrium” towards establishing a time-local arrow of time in the spirit of the Boltzmann H-theorem [15]. While a substantial part of this effort consists in numerical observations, we attempt to trace these finding partially back to the operator growth hypothesis [16] in the second part of the present paper. Throughout this analysis we take the “chaoticity” of a system into account which we measure by the gap ratio [17].
As we find that specifically simple H-theorem type functions emerge (primarily, not exclusively) for chaotic systems, the central quantities developed below add to the limited number of dynamical manifestations of chaos, like, e.g., the correlation hole, described in the literature [18, 19, 20, 21].
The paper is organised as follows. In Sec. II we introduce our construction for arrow of time functions together with the notion of intricacy as a central quantity in the paper at hand. In Sec. III we present numerical data for several pertinent quantum systems and observables. Translating our construction into the setting of the recursion method in Sec. IV we introduce a replacement model to link our findings to the so-called operator growth hypothesis in Sec. V. We conclude by summarizing our work in Sec. VI.
II Construction of arrow of time functions
With the goal to capture the emergent irreversibility of the equilibration process of an observable $\mathcal{O}$ and therefore indicating the direction of time we suggest in the following a scheme to assign the autocorrelation $C(t)=\frac{1}{d}\text{tr}\{\mathcal{O}(t)\mathcal{O}\}$ a function describing an effective deviation to equilibrium. Here $d$ is the dimension of the respective Hilbert space. The scheme iteratively computes a set of $M$ functions $\{\varphi_{j}(t)\}_{j\geq 0}$ starting from $C(t)$ as its germ. Starting with $C(t)=:\varphi_{0}(t)$ the subsequent $\varphi_{j}$ follow via
$\begin{split}\tilde{\varphi}_{j}(t)&:=\partial_{t}\varphi_{j-1}(t)-b_{j-1}\varphi_{j-2}(t),\\ b_{j}^{2}&:=\langle\tilde{\varphi}_{j}(t),\tilde{\varphi}_{j}(t)\rangle,\\ \varphi_{j}(t)&:=-b_{j}^{-1}\tilde{\varphi}_{j}(t),\end{split}$ (1)
with $b_{0}=0$. The inner product $\langle$-,$\rangle$ - , - ⟩ is defined on the basis $\{e_{k}\}=\{\partial_{t}^{k}C(t)\}$ via $\langle e_{n},e_{m}\rangle:=|\partial_{t}^{n+m}C(t)|_{t=0}|$ and constitutes well-defined scalar product on the space spanned by the first $M$ derivatives of $C(t)$. With these $\{\varphi_{j}(t)\}$ we construct
$\alpha^{n}(t):=\sum_{j=0}^{n}\varphi_{j}^{2}(t).$ (2)
We then determine $\mathcal{I}$ as the smallest value for which $\alpha^{\mathcal{I}}$ becomes monotonously decreasing up to a total ascend $\delta$ until the relaxation time $\tau_{\mathcal{I}}$ , the latter being defined by $\alpha^{\mathcal{I}}(t)\leq\delta\quad\forall t\geq\tau_{\mathcal{I}}$. (For the issue of Poincaré recurrences, cf. Ref.  [^1] Strictly speaking such a time $\tau_{\mathcal{I}}$ does not exist for any finite system, since there will always be Poincaré recurrences. These, however occur by magnitudes later than any time scale considered in the paper at hand. Technically $\tau_{\mathcal{I}}$ is determined from $\alpha^{\mathcal{I}}(t)\leq\delta\quad\forall t\geq\tau_{\mathcal{I}}\geq T$ where $T$ is the largest time that we can numerically access. Thus $T$ varies with the considered examples but is always substantially larger than $\tau_{\mathcal{I}}$.) We also refer to the quantity $\mathcal{I}$ as the intricacy of $\alpha$. The value of $\delta$ should be suitably chosen in a sense that is supposed to become clear in the remainder of this paper. For the entirety of the paper at hand we set $\delta=1/200$. For given $C(t),\delta$ a finite $\mathcal{I}$ may or may not exist. We interpret $\alpha^{\mathcal{I}}(t)$ as a deviation-from-equilibrium, e.g., since it limits possible values of $C(t)$ through $\alpha^{\mathcal{I}}(t)\geq C^{2}(t)$. Due to its almost monotonous decrease, we also refer to $\alpha^{\mathcal{I}}(t)$ as an arrow of time function (AOTF). The degree to which the function $\alpha^{\mathcal{I}}(t)$ may be described as time-local depends on $\mathcal{I}$. From the construction one readily infers that $2\mathcal{I}$ temporal derivatives at $t=0$ and $\mathcal{I}$ temporal derivatives at $t$ are required to actually compute $\alpha^{\mathcal{I}}(t)$. Thus, for not too large $\mathcal{I}$, we refer to $\alpha^{\mathcal{I}}(t)$ as time-local with respect to $C(t)$. Physically numerous but in this context somewhat trivial cases emerge for autocorrelation functions $C(t)$ which are themselves monotonously decaying, regardless of the decay time. In this case $\alpha^{0}(t)$ is an AOTF. While this is by no means excluded, this investigation focuses on more complex $C(t)$ that do not decay monotonously but induce AOTFs with finite $\mathcal{I}$. For first examples of AOTFs, see Figs. 1. Note that $C(t)$ in Fig. 1 (bottom) is, unlike all other examples in this paper, not specifically computed from a many-body quantum model, but simply conjectured. Nevertheless a corresponding AOTF may be constructed according to Eqs. (1,2).
To contextualize AOTFs we briefly comment on their comparability to observational entropy [12, 13]. Let $C(t)$ correspond to the (infinite temperature) autocorrelation function of some observable $\mathcal{O}$. Let furthermore $\rho\propto\exp(\beta\mathcal{O})$ with small $\beta$. Then $\alpha^{\mathcal{I}}(t)$ may be computed from the expectation values $\langle\partial_{t}^{n}\mathcal{O}(t)\rangle$ with $0\leq n\leq\mathcal{I}$. Thus, just like observational entropy is based on the knowledge of a specific, case dependent set of expectation values, so is $\alpha^{\mathcal{I}}(t)$. Note however that the $\langle\partial_{t}^{n}\mathcal{O}(t)\rangle$ must not be interpreted as probabilities, as neither $\langle\partial_{t}^{n}\mathcal{O}(t)\rangle>0$ nor $\sum_{n}\langle\partial_{t}^{n}\mathcal{O}(t)\rangle=1$ holds. Thus, the concept of the AOTFs presented here should be considered as complementary to the concept of observational entropy.
Figure 1: Top: Sketch of the functions $\alpha^{n}(t)$ as outlined in (2) for different values for $n$ for the $\sigma_{S}^{x}$ observable in the spin-bath model (16) with $\lambda=0.7$. The index $n$ increases until the construction becomes decreasing up to $\delta$ for some $n=:\mathcal{I}$ (here: $\mathcal{I}=2$). Each $\alpha^{n}(t)$ upper-bounds its predecessor $\alpha^{n-1}(t)$. Inset: The respective autocorrelation function $C(t)$. Bottom: AOTF for the example of robust revivals put forward in Ref. [4]. Concretely we have $C(t)=e^{-5t^{2}}+0.5\left(e^{-5(t-3.5)^{2}}+e^{-5(t+3.5)^{2}}\right)$, see inset. Regardless of the apparent equilibration prior to the revival around $t_{\text{rev}}\sim 3.5$ the AOTF correctly identifies the systems’ eventual approach to equilibrium by giving rise to a monotonous function. Here we find $\mathcal{I}=28$.
III Emergence of arrows of time
In this paper we investigate autocorrelation functions of few-body observables $\mathcal{O}$ in several quantum systems $\mathcal{H}(\sigma)$ which exhibit deviations from the fully chaotic regime moderated by specific system parameters $\sigma$. We quantify the notion of chaoticity by means of the mean gap ratio $\langle r\rangle$ of adjacent level spacings as a standard indicator for quantum chaos [17]. Concretely by
$\langle r\rangle=\frac{1}{d}\sum_{k}\frac{\min(s_{k},s_{k+1})}{\max(s_{k},s_{k+1})},$ (3)
where $s_{k}=\epsilon_{k+1}-\epsilon_{k}$ with the $\epsilon_{k}$ as the eigenenergies of the considered system. Exploiting respective symmetries, we truncate the system to a specific symmetry separated subspace. For fully chaotic systems, the mean gap ratio is numerically determined as $\langle r\rangle_{\text{GOE}}\approx 0.53$ by means of large random matrices from a Gaussian orthogonal ensemble (GOE). Conversely, for integrable systems exhibiting a Poissonian spectrum, $\langle r\rangle_{P}\approx 0.386$ [17].
Whilst tracking the respective gap ratio along $\sigma$ we also compute the autocorrelation $C(t)$ for different system parameters. For these autocorrelation functions we determine $\mathcal{I}$ as the intricacy of the arrow of time function as described in Sec. II.
For details on the numerical method applied in this work we refer to Appendix A.
III.1 Models and observables
Two-chain model: As a first model we consider two (mixed-field) Ising chains with local coupling of the respective $\sigma^{z}$-terms moderated by the interaction strength $\lambda$ and with periodic boundary conditions. The Hamiltonian reads
$\mathcal{H}(\lambda)$ $=\mathcal{H}_{1}+\lambda\mathcal{H}_{I}+\mathcal{H}_{2},$ (4)
with the Ising Hamiltonians
$\mathcal{H}_{a}=\sum_{k}h_{a,k},$ (5)
with $a=1,2$, local constituents
$\begin{split}h_{a,k}&=J\sigma_{a,k}^{z}\sigma_{a,k+1}^{z}+\frac{B_{x}}{2}\left(\sigma_{a,k}^{x}+\sigma_{a,k+1}^{x}\right)\\ &+\frac{B_{z}}{2}\left(\sigma_{a,k}^{z}+\sigma_{a,k+1}^{z}\right),\end{split}$ (6)
and the interaction
$\mathcal{H}_{I}=\sum_{k}\sigma_{1,k}^{z}\sigma_{2,k}^{z}.$ (7)
We fix the parameters of the model as $(J,B_{x},B_{z})=(1,1,0.5)$ and vary the interaction strength $\lambda$. For the computation we consider a total size $L=18$.
As the model mimics the pertinent scenario of the equilibration of two macroscopic bodies in contact, we consequently consider the energy imbalance
$\mathcal{O}=\mathcal{H}_{1}-\mathcal{H}_{2}.$ (8)
For low $\lambda$ an (almost) exponentially decaying $C(t)$ results, see Fig. 2. This is reminiscent of the decay of a temperature difference between to macroscopic bodies in mutual thermal contact, but insulated otherwise. The application of a standard projection operator technique in the Markovian, weak coupling limit would find this result [7]. The intricacy is $\mathcal{I}=0$. However, as we increase $\lambda$ we leave the Markovian regime and at $\lambda\sim 2$ $C(t)$ is no longer monotonously decaying but exhibits a more intricate form, see Fig. 2. Nevertheless the intricacy is moderate , $\mathcal{I}=1$ and $\langle r\rangle$ is still close the GOE value, cf. Fig. 3. Only if $\lambda$ is increased even further towards $\lambda\sim 2.5$ the gap ratio $\langle r\rangle$ sinks below the GOE value (since the model approaches a situation of decoupled dimers) and the intricacy substantially increases to $\mathcal{I}=86$.
Figure 2: Autocorrelation functions $C(t)$ (left) and their respective squares $C^{2}(t)$ and AOTFs $\alpha^{\mathcal{I}}$ (right) for energy imbalance in the two-chain-model (4) for different interaction strengths $\lambda$. Top: At $\lambda=0.7$ (chaotic regime, see Fig. 3) $C(t)$ itself is monotonously decreasing and hence  $\alpha^{0}$ is an AOTF . Centre: At $\lambda=2.2$ (chaotic regime, see Fig. 3) $C(t)$ no longer decreases monotonously. However, $\alpha^{1}$ does, thus two derivatives w.r.t. time are required to construct an AOTF . Bottom: In the nonchaotic regime at $\lambda=2.7$ (see Fig. 3) a total of 260 derivatives w.r.t. time are required to construct the AOTF $\alpha^{130}$.
In the direction of even larger $\lambda$, i.e. leaving the fully chaotic regime, we observe further growth of $\mathcal{I}$, see Fig. 3. For this model we can rule out arrows of time for $\mathcal{I}\leq 200$ for $\lambda\gtrsim 2.8$. Conversely, in the chaotic regime we find arrows of time with low intricacy giving rise to a time-local description of the approach to equilibrium as only a low finite numbers of derivatives of the autocorrelation function $C(t)$ at times 00 and $t$ are sufficient to construct an AOTF.
Figure 3: AOTF intricacy $\mathcal{I}$ and gap ratio $\langle r\rangle$ for different interaction strengths $\lambda$ in the site-wise coupled two-chain-model with $L=18$. The black line indicates the chaotic benchmark $\langle r\rangle_{\text{GOE}}=0.53$ and the shaded area $1\%$ deviation around it. The red arrow indicates that for larger $\lambda$ our numerical method suffices to determine resp. $\mathcal{I}$-values to exceed 200.
XXZ chain: Next, we consider an XXZ-type Heisenberg chain with length $L=20$ and periodic boundary conditions, modelled by the Hamiltonian
$\mathcal{H}(\Delta)=J\sum_{k=1}^{L}s_{k}^{x}s_{k+1}^{x}+s_{k}^{y}s_{k+1}^{y}+\Delta s_{k}^{z}s_{k+1}^{z}+\Delta^{\prime}s_{k}^{z}s_{k+2}^{z},$ (9)
in which we set $(J,\Delta^{\prime})=(1,0.5)$ and where $s_{k}^{a}=\sigma_{k}^{a}/2$ denote the spin-matrices. For the entirety of this paper we set $\hbar=1$. We vary the nearest-neighbour interaction term $\Delta$. Introducing the next-nearest interaction $\Delta^{\prime}$ breaks Bethe-Ansatz integrability [23, 24]. However, for $\Delta\gg\Delta^{\prime}$ the system again approaches the integrable regime, see Fig. 5. For any isotropies $\Delta,\Delta^{\prime}$ the total magnetization is conserved [23], motivating the study of the respective current
$\mathcal{O}=\mathcal{J}_{S}=\sum_{k}s^{x}_{k}s^{y}_{k+1}-s^{y}_{k}s^{x}_{k+1}$ (10)
as an observable of interest. Depictions of respective AOTFs are given in Fig. 4. Again, we find that for the chaotic system the autocorrelation functions allow for simple AOTFs with $\mathcal{I}\approx 5$. The intricacy $\mathcal{I}$ does grow to large values of $\mathcal{I}>40$ but this happens deeper in the nonchaotic regime, at $\Delta\approx 10$, see Fig. 5.
Figure 4: Autocorrelation functions $C(t)$ (left) and their respective squares $C^{2}(t)$ and AOTFs $\alpha^{\mathcal{I}}$ (right) for magnetisation current in the XXZ model (9) for different nearest-neighbour interactions $\Delta$. Top: At $\Delta=3$ (chaotic regime, see Fig. 5) $C(t)$ does not decay monotonously but the AOTF $\alpha^{3}(t)$ does. Bottom: Further away from the chaotic regime, at $\Delta=10$, an AOTF is reached only at $\mathcal{I}=47$, i.e. $\alpha^{47}(t)$ decays monotonously.
Figure 5: AOTF intricacy $\mathcal{I}$ and gap ratio $\langle r\rangle$ for different nearest-neighbour couplings $\Delta$ in the XXZ Heisenberg chain with $L=20$. The black line indicates the chaotic benchmark $\langle r\rangle_{\text{GOE}}=0.53$ and the shaded area $1\%$ deviation around it.
Ising chain: As a third model we turn to a transverse Ising chain with length $L=18$ with a tilted magnetic field as very generic example for a chaotic quantum model. Its Hamiltonian is given by
$\mathcal{H}(B_{x})$ $=\sum_{k}h_{k},$ (11)
with the local energies
$\begin{split}h_{k}&=J\sigma_{k}^{x}\sigma_{k+1}^{x}+\frac{B_{z}}{2}\left(\sigma_{k}^{z}+\sigma_{k+1}^{z}\right)+\frac{B_{x}}{2}\left(\sigma_{k}^{x}+\sigma_{k+1}^{x}\right).\end{split}$ (12)
We choose the parameter set as $(J,B_{z})=(1,-1.05)$ and employ periodic boundary conditions. We vary $B_{x}$ breaking integrability as for $B_{x}=0$ the system is integrable whereas for $B_{x}\neq 0$ the system enters the chaotic regime. We investigate the arrows of time of several kinds of physically interesting observables. First we consider
$\mathcal{O}_{1}\propto\sum_{k}\cos\left(\pi k\right)h_{k}$ (13)
representing the fast-mode of the energy density-wave operator. This operator shows a quick relaxation.
Moreover we also examine the 2-local observable
$\mathcal{O}_{2}\propto\sum_{k}1.05\ \sigma_{k}^{x}\sigma_{k+1}^{x}+\sigma_{k}^{z}.$ (14)
In Fig. 6 we depicted exemplary autocorrelation functions of $\mathcal{O}_{2}$ as well as respective AOTFs for particular magnetic fields $B_{x}$ in the chaotic regime as well as close to the integrable point.
Lastly, we include to our investigation the energy current
$\mathcal{O}_{3}=\mathcal{J}_{E}=B_{x}\sum_{k}\sigma_{k}^{y}(\sigma_{k+1}^{z}-\sigma_{k-1}^{z}).$ (15)
Before delving into the analysis of our findings for this model, some remarks are in order. In the integrable Ising chain with $B_{x}=0$ the current $\mathcal{J}_{E}$ is an integral of motion. Consequently the respective autocorrelation function never relaxes and no intricacy $\mathcal{I}$ can be inferred subject to the above scheme (1). At small magnetic fields $B_{x}$ a possible relaxation is beyond the time scale we consider, hence we do not include these points here.
Whereas for $B_{x}=0$ the model (11) is integrable, introducing a non-zero tilted field $B_{x}$ breaks integrability and the system becomes chaotic, see Fig. 7. Again, deep in the chaotic regime we find simple, low-intricacy AOTFs for all considered observables. Approaching the integrable point as $B_{x}\searrow 0$, for the observables $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ the intricacies $\mathcal{I}$ grow, eventually exceeding the scope of our numerical analysis at $B_{x}\lesssim 0.125$ indicating intricacies $\mathcal{I}>100$. For the energy current $\mathcal{O}_{3}$ the intricacies are low for every considered magnetic field $B_{x}$.
Figure 6: Autocorrelation functions $C(t)$ (left) and their respective squares $C^{2}(t)$ and AOTFs $\alpha^{\mathcal{I}}$ (right) for the orthogonal observable $\mathcal{O}_{2}$ in the tilted field Ising chain (11) for different magnetic fields $B_{x}$. Top: At $B_{x}=0.2$ (chaotic regime, see Fig. 7) the autocorrelation function allows for an AOTF with intricacy $R=7$. Bottom: Approaching the fully integrable regime at $B_{x}=0$ (see Fig. 7), we find for $B_{x}=0.125$ an AOTF with $\mathcal{I}=94$.
Figure 7: AOTF intricacy $\mathcal{I}$ and gap ratio $\langle r\rangle$ for magnetic fields $B_{x}$ in the tilted field Ising chain with $L=18$. The black line indicates the chaotic benchmark $\langle r\rangle_{\text{GOE}}=0.53$ and the shaded area $1\%$ deviation around it. The arrows indicate that for smaller $B_{x}$ our numerical method suffices to determine resp. $\mathcal{I}$-values to exceed 100.
Spin-bath model: Lastly we investigate the case of a single spin-1/2 coupled to an Ising spin bath of length $L=14$ with periodic boundary conditions. This system is described by the Hamiltonian
$\mathcal{H}(\lambda)=\mathcal{H}_{S}+\lambda\mathcal{H}_{I}+\mathcal{H}_{B},$ (16)
where $\lambda$ moderates the interaction of the spin and the bath via a coupling of the respective $x$-components. Concretely, we have
$\mathcal{H}_{S}$ $=\sigma_{S}^{z}$ (17)
$\mathcal{H}_{I}$ $=\sigma_{S}^{x}\otimes\sigma_{1}^{x}$ (18)
$\mathcal{H}_{B}$ $=J\sum_{j}\sigma_{j}^{z}\sigma_{j+1}^{z}+B_{x}\sigma_{j}^{x}+B_{z}\sigma_{j}^{z},$ (19)
where we choose the parameters as $(\omega,J,B_{x},B_{z})=(1,1,1,0.5)$. Varying the coupling strength $\lambda$ we compute the autocorrelation for the $x$ and $z-$component of coupled spin respectively, i.e. $\mathcal{O}^{(x)}=\sigma_{S}^{x}$ and $\mathcal{O}^{(z)}=\sigma_{S}^{z}$. See Figs. 1 and 8 for exemplary AOTFs.
Figure 8: Autocorrelation functions $C(t)$ (left) and their respective squares $C^{2}(t)$ and AOTFs $\alpha^{\mathcal{I}}$ (right) for the observable $\mathcal{O}_{S}^{z}$ in the spin bath model (16) for different couplings $\lambda$. Top: At small coupling $\lambda$, here: $\lambda=2$ (chaotic regime, see Fig. 9) we find an AOTF with intricacy $\mathcal{I}=10$. Bottom: Approaching the chaotic regime again at $\lambda=12$, the intricacy of the respective AOTF drops to $\mathcal{I}=48$ again falling within our numerical scope.
The spin-bath model (16) exhibits a deviation from full chaoticity at intermediate spin-bath coupling $4\lesssim\lambda\lesssim 11$, cf. Fig. 9.
At small coupling $\lambda$ the intricacies $\mathcal{I}$ for the AOTFs of $\sigma_{S}^{z}$ start off at $\mathcal{I}=0$ as at first the autocorrelation functions are monotonous themselves. With increasing coupling strength, the intricacies grow until $\mathcal{I}=19$ for $\lambda=2.5$. As the system approaches the nonchaotic regime the intricacies exceed $\mathcal{I}=100$, leaving the scope of our numerical method. Once the system approaches the chaotic regime again at $\lambda\sim 11$ we again infer intricacies $\mathcal{I}<100$ starting at $\lambda=12$ where $\mathcal{I}=48$. These then decrease further as the system moves away from the nonchaotic regime, eventually giving rise to an AOTF with $\mathcal{I}=5$ at $\lambda=20$.
In the case of the observable $\sigma_{S}^{x}$ the picture is different. The intricacies do not seem to be affected by the systems’ transition from the chaotic to an nonchaotic regime. With an exception at $\lambda=2$ where our construction finds the respective autocorrelation function to be suitably monotonous, i.e. $\mathcal{I}=0$, we find AOTFs with an intricacy $\mathcal{I}=2$ for any other considered coupling strength $\lambda\leq 8.5$. Entering a strong coupling range at $\lambda\gtrsim 9$ however, the functions $C^{2}(t)$ do no longer decay within a reasonable period in the sense we require to infer an equilibration time $\tau$. Consequently we leave them out of further discussions.
Having presented all numerical evidence, we now aim at distilling the common principle that is present in all these examples: Whenever an autocorrelation function of a few body observable exhibits high intricacy, e.g. $\mathcal{I}>20$ (for the $\delta$ chosen here) the corresponding system either belongs to the nonchaotic regime (according to the gap ratio) or a nonchaotic regime is close by with respect to variation of some model parameter. Note that the converse does not hold: low intricacy dynamics may emerge in nonchaotic regimes, see e.g. $\sigma_{S}^{z}$ in the spin-bath model. This principle is our main numerical result.
The implication of this principle may be loosely phrased as: Deep in the chaotic regime dynamics are simple and time-local arrows of time may be readily constructed. Intricate dynamics require at least the closeness of a nonchaotic regime.
Figure 9: AOTF intricacy $\mathcal{I}$ and gap ratio $\langle r\rangle$ for different coupling strengths $\lambda$ in the model of a single spin coupled to an Ising chain with $L=14$. The black line indicates the chaotic benchmark $\langle r\rangle_{\text{GOE}}=0.53$ and the shaded area $1\%$ deviation around it. The arrows indicate that within the parameter range $3\lesssim\lambda\lesssim 11$ the intricacies exceed $\mathcal{I}=100$.
IV Arrows of time in the Lanczos picture
In the second part of the paper at hand we wish to put our numerical findings into the context of well-established concepts such as the recursion method and Lanczos coefficients [25, 16]. To this end we briefly introduce pertinent concepts, translate our construction into this setting and expand on relevant quantities.
IV.1 The Recursion Method and Lanczos coefficients
Given some Hamiltonian $\mathcal{H}$ and Hermitian operator $\mathcal{O}$, we are interested in autocorrelation functions
$C(t)=\text{tr}\{\mathcal{O}(t)\mathcal{O}\}/d,$
where we denote with $\mathcal{O}(t)$ the Heisenberg time evolution $U(t)^{\dagger}\mathcal{O}U(t)$ for $U(t)=\exp(-i\mathcal{H}t)$ and $d$ the dimension of the Hilbert space. Working in operator space, also Liouville space, we denote the vectors by $|\mathcal{O})$. Equipped with the Hilbert-Schmidt scalar product $(A|B):=\text{tr}\{A^{\dagger}B\}/\text{tr}\{\mathbf{1}\}$ a norm is induced via $||\mathcal{O}||=\sqrt{(\mathcal{O}|\mathcal{O})}$. The Liouville superoperator $\mathcal{L}(-):=\left[\mathcal{H},-\right]$, also Liouvillian, here implements the time evolution as $|\mathcal{O}(t))=\exp\left(i\mathcal{L}t\right)|\mathcal{O})$. With the Liouvillian it follows that for the autocorrelation we have $C(t)=(\mathcal{O}|\exp(i\mathcal{L}t)|\mathcal{O})$.
The Lanczos algorithm now orthogonalizes the set $\{\mathcal{L}^{n}|\mathcal{O})\}$ as follows. Starting from $|\mathcal{O}_{0})=:|\mathcal{O})$ the algorithm iteratively computes the Krylov basis $\{|\mathcal{O}_{n})\}$ as
$\begin{split}|\tilde{\mathcal{O}}_{n})&=\mathcal{L}|\mathcal{O}_{n-1})-b_{n-1}|\mathcal{O}_{n-2})\\ b_{n}&=(\tilde{\mathcal{O}}_{n}|\tilde{\mathcal{O}}_{n})^{1/2}\\ |\mathcal{O}_{n})&=b_{n}^{-1}|\tilde{\mathcal{O}}_{n}),\end{split}$ (20)
where $b_{0}=0$. Upon defining the functions
$\varphi_{n}(t):=i^{-n}(\mathcal{O}_{n}|\mathcal{O}(t))$ (21)
the connection to the scheme (1) becomes obvious. The time derivatives of the autocorrelation function in the AOTF construction correspond to the repeated action of the Liouvillian onto the respective observable. Thus the functions $\varphi_{n}(t)$ are equal to the overlap of the $n^{\text{th}}$ Krylov vector with $\mathcal{O}(t)$ as explained in (21). The quantities $b_{n}$ are by construction real and positive and are referred to as Lanczos coefficients.
With the definition (21) comes a frequently adapted interpretation of the Lanczos coefficients as hopping amplitudes of a tight-binding model. The entries of the wave function on this semi-infinite chain are given by (21). For a sketch see Fig. 10. The wave vector satisfies a discrete Schrödinger equation
$\partial_{t}\varphi_{n}(t)=b_{n}\varphi_{n-1}(t)-b_{n+1}\varphi_{n+1}(t)\quad\text{with}\ n\geq 0,$ (22)
where we set $\varphi_{-1}=0$ by convention and with initial condition $\varphi_{n}(0)=\delta_{n,0}$. This may be cast into the form of a matrix equation
$\dot{\boldsymbol{\varphi}}=\textbf{L}.\boldsymbol{\varphi},$ (23)
where the matrix L is given by
L $=\begin{pmatrix}0&-b_{1}&0&0&\dots\\ b_{1}&0&-b_{2}&0&\dots\\ 0&b_{2}&0&-b_{3}&\dots\\ 0&0&b_{3}&0&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\\ \end{pmatrix},$ (24)
and where $\boldsymbol{\varphi}(t)=(\varphi_{0}(t),\varphi_{1}(t),\dots)$.
Figure 10: Sketch of the wave vector (21) for different instances of time $t$ at the example of fast-mode observable $\mathcal{O}_{1}$ of the energy density in the tilted field Ising chain with $B_{x}=0.6$.
IV.2 Currents and probabilities
The analogy of the tight-binding model gives rise to the interpretation of the entries $\varphi_{n}(t)$ of the wave vector as amplitudes for a particle being at some site $n$ with probability $\varphi_{n}^{2}(t)$ as we have
$\sum_{n=0}^{\infty}\varphi_{n}^{2}(t)=1.$ (25)
We then define a probability operator $\textbf{P}_{k}=\mathbf{1}_{k}\oplus\mathbf{0}$ whose expectation value
$\langle\textbf{P}_{k}(t)\rangle=\langle\boldsymbol{\varphi}(t)|\textbf{P}_{k}|\boldsymbol{\varphi}(t)\rangle=\sum_{j=0}^{k}\varphi_{j}^{2}(t)$ (26)
describes the probability to be within the first $k$ sites. From this we can read off that
$\alpha^{\mathcal{I}}(t)=\langle\textbf{P}_{\mathcal{I}}(t)\rangle.$ (27)
The dynamics of (26) is determined by the Liouvillian L as
$\langle\dot{\textbf{P}}_{k}\rangle=\langle\left[\textbf{P}_{k},\textbf{L}\right]\rangle=-2b_{k}\varphi_{k}(t)\varphi_{k+1}(t)=:-\langle\textbf{J}_{k}(t)\rangle,$ (28)
where we define $\textbf{J}_{k}$ as the associated probability current along the $k$-th site.
Hence, an AOTF with intricacy $\mathcal{I}$ is equivalent to the probability of the particle on the semi-infinite chain to be within the first $\mathcal{I}$ sites. This probability is (suitably) monotonously decreasing, i.e. there are no (few) backward currents of probability.
Moreover, from the comparison of $\alpha^{\mathcal{I}}$ with the dynamics of the wave vector $\boldsymbol{\varphi}(t)$ we can conclude that whenever $\boldsymbol{\varphi}(t)$ has a zero, i.e. two adjacent entries have opposite signs, the AOTF must not be monotonous by (28). Henceforth we argue that by this it is reasonable to allow for small errors $\delta$ from a monotonous behaviour as by continuity is not to be expected that such zeros only occur locally but travel through the chain.
V structures in the Lanczos coefficients and the arrow of time
Having translated the AOTF construction put forward in Sec. II into the Lanczos picture in Sec. IV we further want to exploit this correspondence by bringing the dynamics of the AOTF into relation with structures in the Lanczos coefficients. To this we motivate an effective replacement model for the dynamics of the wave vector $\boldsymbol{\varphi}(t)$ and discuss its connection to the so-called operator growth hypothesis [16].
V.1 Replacement model
In order to provide a more intuitive picture of the dynamics of the wave vector $\boldsymbol{\varphi}(t)$ in the Lanczos picture, we aim at mapping Eq. (22) onto a replacement model. We emphasize that the profit of this mapping consists more in the provision of a simple comparable model for many sets of $b_{n}$, rather than its strict applicability for all possible sets of $b_{n}$. Discussing the latter in detail would exceed the scope of the paper at hand.
To this end we consider the dynamics of a particle with mass $m=1/2$ in one dimension under the influence of a potential $V=-B^{2}$. With $\psi(x,t)$ as the corresponding wave function the dynamics are given by
$i\partial_{t}\psi(x,t)=\left[-\partial_{x}^{2}-B^{2}(x)\right]\psi(x,t).$ (29)
Upon a gauge transformation
$\psi(x,t)=\exp\left(i\phi(x)\right)\vartheta(x,t),$ (30)
where $\phi^{\prime}(x)=B(x)$ one can readily infer that for the dynamics of $\vartheta(x,t)$ we have
$i\partial_{t}\vartheta(x,t)=-\left[2iB(x)\partial_{x}+iB^{\prime}(x)+\partial_{x}^{2}\right]\vartheta(x,t),$ (31)
where $B^{\prime}(x)=\text{d}B(x)/\text{d}x$. Before proceeding the following should be noted: A lowest order WKB-approximation to the wave function of an outgoing particle with energy eigenvalue $E=0$ for the Hamiltonian in Eq. (29) reads:
[…]
Applying the gauge transformation Eq. (30) to $\psi_{\text{WKB}}$ precisely removes the possibly quickly oscillating phase factor, yielding $\vartheta_{\text{WKB}}\propto 1/\sqrt{B(x)}$. At eigenenergies close to $E=0$ the phase factor will not be removed completely but transformed to oscillate slowly. Thus the following may be expected: For potential landscapes $V=-B^{2}(x)$ to which the WKB-approximation applies (no turning points of the classical particle, etc.) and for initial states restricted to low energies, the curvature of $\vartheta(x,t)$ will be and remain low, i.e. $\partial_{x}^{2}\vartheta(x,t)\approx 0$ and the corresponding term in Eq. (31) may be neglected.
Turning back to the dynamics of the wave function $\boldsymbol{\varphi}(t)$ in the Lanczos picture, we consider the continuum limit of (22), i.e. where $b_{n}\rightarrow b(x),\ \varphi_{n}(t)\rightarrow\varphi(x,t)$, cf. [16, 26]. Note that whereas the Lanczos coefficient $b_{n}$ is “located” between the sites $n-1$ and $n$, the function $b(x)$ takes values at position $x$. Hence $b_{n}\ \widehat{=}\ b(x-\frac{1}{2})$. Up to first order of $b(x)$ and $\varphi(x,t)$ in $x$ we find for (22) that
$\partial_{t}\varphi(x,t)=-\left[2b(x)\partial_{x}+b^{\prime}(x)\right]\varphi(x,t),$ (33)
[…]
$\frac{\text{d}^{2}\langle\hat{x}\rangle}{\text{d}t^{2}}=-2\langle\nabla V(\hat{x})\rangle.$ (34)
Correspondingly, for the respective quantity in the Lanczos picture we compute
$\frac{\text{d}^{2}\langle\textbf{n}\rangle}{\text{d}t^{2}}$ $=\langle\left[\left[\textbf{n},\textbf{L}\right],\textbf{L}\right]\rangle=2\sum_{j}\left(b_{j+1}^{2}-b_{j}^{2}\right)\varphi_{j}^{2},$ (35)
where $\textbf{n}=\text{diag}(1,2,\dots)$, again suggesting the form of the potential to be $V(x)=-b^{2}(x)$.
Having motivated the analogy of the Lanczos dynamics to a particle exposed to a potential $-b^{2}(x)$ we seek to conclude on properties of $\boldsymbol{\varphi}(t)$ based on the simple replacement model.
In order for a particle to propagate and to not be trapped, the potential $V$ necessarily needs to fall off, i.e. the Lanczos coefficients need to ascend. Furthermore, on qualitative grounds we argue that for a steady “forward-oriented” motion of the particle, in the sense of no back-transport, the potential additionally needs to exhibit a certain smoothness in order to not give rise to scattering or localization. For this second property we deliberately remain unspecific as we only aim to further stress the suitability of the replacement model rather than to quantify notions of smoothness.
V.2 Operator Growth Hypothesis
The operator growth hypothesis (OGH) is a statement on the growth of Lanczos coefficients in infinite-dimensional chaotic many-body systems [16]. It postulates that in such systems the Lanczos coefficients for local observables asymptotically attain linear growth, i.e.
$b_{n}=\left\{\begin{array}[]{ll}A\frac{n}{\ln n}+o\left(\frac{n}{\ln n}\right)&d=1,\\ \alpha n+\gamma+o(1)&d>1.\end{array}\right.$ (38)
For numerical purposes however, one is limited to finite systems. In such systems of size $L$, the first $L/2$ Lanczos coefficients $b_{n}$ coincide with those of the infinite system. Several studies suggest that already the first several Lanczos coefficients feature a linear trend [16, 28, 29, 30, 31]. However whether this precise behaviour persists forever clearly lies beyond the scope of any numerical investigation and is no part of the OGH.
V.3 Lanczos Coefficients and Intricacies
The analogy between the Lanczos picture and the replacement model predicts that for a steady outward motion of $\boldsymbol{\varphi}(t)$ the Lanczos coefficients need to grow with larger $n$. Moreover the OGH puts forward that in chaotic quantum systems the Lanczos coefficients of local few-body observables eventually attain linear growth. Building up on this connection between Lanczos coefficients and the dynamics of the wave vector $\boldsymbol{\varphi}(t)$ we seek to bridge these structural arguments to the intricacies $\mathcal{I}$ of the respective AOTFs. As pointed out in Sec. IV.2, an AOTF with intricacy $\mathcal{I}$ can be thought of in the Lanczos picture as the (almost) monotonously decreasing probability of the particle to be located within the first $\mathcal{I}$ sites and is thus directly correlated to the dynamics of $\boldsymbol{\varphi}(t)$.
Straight-forwardly we determine the position $P$ after which the Lanczos coefficients have terminally surpassed some multiple of $b_{1}$. Notice that rescaling the set of Lanczos coefficients changes only the time-scale of the associated dynamics. We consider the value $1.5\ b_{1}$, see Fig. 11. Note that we consider finite systems. Hence the OGH only pertains to the first $L/2$ Lanczos coefficients, where $L$ is the size of the system. The growth of the Lanczos coefficients does not prevail substantially longer due to the finiteness of the system, e.g. cf. Fig. 11 $g)$ around $n\sim 25$. In Fig. 11 we only depict as many Lanczos coefficients as necessary to depict every $P$ in order to avoid cluttering.
Figure 11: First several Lanczos coefficients for the models and observables considered in Sec. III. In each plot the Lanczos coefficients are shifted vertically for presentation purposes. The dashed line indicates the position of the first Lanczos coefficient. To emphasise the growth, we included $b_{0}=0$ for each set $\{b_{n}\}$. The black dots correspond to the points $P$ where the Lanczos coefficients have terminally surpassed $1.5\ b_{1}$. In cases where $\mathcal{I}>100$ we colour the respective $b_{n}$ black regardless whether we are able to infer an $\mathcal{I}$ or not. a) Ising ladder, b) XXZ chain, c)-e) $\mathcal{O}_{1}-\mathcal{O}_{3}$ in the tilted field Ising chain respectively, f),g) $\mathcal{\sigma}_{S}^{x},\mathcal{\sigma}_{S}^{z}$ in the spin-bath model. For the coupling strength $\lambda=20$, the Lanczos coefficients for $\sigma_{S}^{z}$ do not terminally surpass 1.5 $b_{1}$. Therefore we cannot infer $P$ in this case.
We find for each pair $(\mathcal{H},\mathcal{O})$ that whenever the growth of the Lanczos coefficients sets in early (as quantified by $P$), the corresponding intricacy $\mathcal{I}$ is low. The absolute values for $\mathcal{I}$ however are not comparable between different systems and observables. In the derivation for the replacement dynamics the structure of the Lanczos coefficients only enters up to first order. For the two-chain model (4), the XXZ chain (9) as well as for the tilted field Ising model (11) the Lanczos coefficients mostly do not exhibit strong non-linear behaviour, see Fig. 11 $a)$-$e)$. In these cases $P$ and the intricacies $\mathcal{I}$ correlate. This observation also holds true for the spin-bath model (16) for small coupling $\lambda$, as depicted in Fig. 11 $f)$, $g).$ However starting at intermediate coupling strength $\lambda\sim 3$ the Lanczos coefficients exhibit a non-linear, staggered structure, already for the infinite-system $b_{n}$. Whereas at first the correlation between $P$ and $\mathcal{I}$ is visible for small spin-bath coupling strengths $\lambda$ for $\mathcal{O}=\sigma_{S}^{z}$, the analogy fails in the strong-coupling regime. At $\lambda=20$ the intricacy becomes $\mathcal{I}=5$, although $P$ is largest for this interaction strength, exceeding the range of Lanczos coefficients considered for this example. However, as already pointed out above, the dynamics for the replacement model only takes the structure of the $b_{n}$ up to first order into account.
Nevertheless, we find that within its range of applicability the reasoning based on the replacement model strongly hints on a connection between the intricacies $\mathcal{I}$ and growth of the respective Lanczos coefficients. Furthermore we expect that whenever the OGH holds true in the sense that the linear growth not only prevails asymptotically but also sets in conceivably early (as observed in several studies), the intricacies of AOTFs are small.
VI conclusion and remarks
In this work we proposed a construction for monotonously decreasing arrow of time functions (AOTFs) for autocorrelation functions of local observables. The construction of the AOTFs requires $2\mathcal{I}$ temporal derivatives of the respective autocorrelation function itself. Hence the intricacy $\mathcal{I}$ characterises the extend to which an AOTF may be considered to be time-local with respect to the autocorrelation function.
We numerically investigated the AOTFs of several few-body observables in pertinent quantum models. In these we found that whenever the system is fully chaotic (subject to the gap ratio) the intricacies $\mathcal{I}$ are low, allowing for a time-local description of the systems’ approach to equilibrium. Conversely, only when the system is nonchaotic or close to a nonchaotic regime (with respect to some model parameter) the dynamics may become highly intricate. However, integrability itself (or closeness to it) is no sufficient condition for high-intricacy dynamics.
Furthermore we translated our findings into the setting of Lanczos coefficients and the recursion method. In this picture the AOTF is equivalent to the probability of a particle to be located within the first $\mathcal{I}$ sites in a tight-binding model, where the Lanczos coefficients $b_{n}$ take the role of the hopping amplitudes.
By linearising the continuum limit of the equations of motion in the Lanczos picture we introduced a replacement model providing an analogy to the motion of particle in a potential given by the negative square of the Lanczos coefficients $b_{n}$. In this analogy the motion of the particle in the replacement model represents the flux of probability in the tight-binding model. Consequently if the motion in the replacement model is predominantly outbound, there are only few backward currents of probability in the Lanczos picture which in turn gives rise to an arrow of time. In order for this to occur the potential needs to decrease, which in turn requires growing Lanczos coefficients. This is to be compared to the operator growth hypothesis which predicts growing Lanczos coefficients for chaotic systems. We find in most of the cases the linearisation to be appropriate for the first several Lanczos coefficients. In these cases the growth of the $b_{n}$ and the intricacies correlate, linking our our numerical findings to the operator growth hypothesis.
Acknowledgements
We thank Christian Bartsch, Robin Heveling and Mats H. Lamann for fruitful discussions. JW thanks Heiko G. Menzler for discussions on the numerical method. This work has been funded by the Deutsche Forschungsgemeinschaft (DFG), under Grant No. 531128043, as well as under Grant No. 397107022, No. 397067869, and No. 397082825 within the DFG Research Unit FOR 2692, under Grant No. 355031190.
Appendix A Numerical method
For the numerics in Sec. III we determine the functions $\varphi_{j}(t)$ defined in (20) as entries of the wave vector $\boldsymbol{\varphi}$ via (21). To ensure orthogonality between the vectors $|\mathcal{O}_{n})$ in the Krylov basis we perform partial re-orthogonalisation [32, 33]. Additionally we numerically verify that $\max_{n}|\partial_{t}\varphi_{n}-b_{n}\varphi_{n-1}+b_{n+1}\varphi_{n+1}|\leq 10^{-8}$, indicating the accuracy of our result.
Appendix B Derivation of (33)
Linearising both $\varphi(x)$ and $b(x)$ in $x$ we have
$\varphi_{n\pm 1}(t)$ $\rightarrow\varphi(x)\pm\varphi^{\prime}(x),$
$b_{n}$ $\rightarrow b(x)-\frac{1}{2}b^{\prime}(x),$
$b_{n+1}$ $\rightarrow b(x)+\frac{1}{2}b^{\prime}(x).$
With this we one can rewrite (22) to find (33).